# A Compositional Framework for On-the-Fly ltlf Synthesis

### v1

@@ 1 Introduction
We implemented our synthesis framework in a prototype tool called Cosy. To evaluate the efficiency of the proposed composition techniques, we conducted an empirical evaluation by comparing Cosy with state-of-the-art ltlf synthesizers. Experimental results show that our compositional synthesis framework solves a notable number of instances that other synthesizers cannot handle, demonstrating that it effectively combines the merits of both backward and forward synthesis approaches for ltlf specifications. A detailed comparison of the two composition variants shows that, while the incremental variant solves slightly more instances, both variants have their strengths and advantages.
We continue in the next section with preliminaries. Sections 3 and 4 present compositional on-the-fly ltlf synthesis from the theoretical foundations and implementation aspects, respectively. Section 5 shows the experimental results. Finally, Section 6 summarizes the contributions and discusses future work.
@@ 2 Preliminaries
Definition 1 (ltlf Realizability and Synthesis).
An ltlf synthesis specification $\left(\varphi,\mathcal{X},\mathcal{Y}\right)$ is realizable, iff there exists an agent strategy $g:(2^{\mathcal{X}})^{*}\to 2^{\mathcal{Y}}$ that is winning, such that for every infinite sequence $\lambda=X_{0}X_{1}\cdots\in(2^{\mathcal{X}})^{\omega}$ of propositional interpretations over $\mathcal{X}$, i.e., every possible environment behaviours, there is $k\geq 0$ such that $\rho\models\varphi$ holds, where $\rho=(X_{0}\cup g(\epsilon))(X_{1}\cup g(X_{0}))\cdots(X_{k}\cup g(X_{0}\cdots X_{k-1}))$. The ltlf realizability problem is to determine whether an ltlf specification is realizable, and the ltlf synthesis problem for a realizable specification is to compute a winning strategy. An ltlf specification is unrealizable if it is not realizable.
@@ 3 Theoretical Foundations
To formulate the composition of strategies, we utilize the agent-winning region in dfa games. Intuitively, the agent-winning region, consisting of all agent-winning states, captures all possible ways the agent can win the game, regardless of how the environment behaves. Consequently, it represents all the information required from each sub-specification to realize the original specification. Notably, the environment-winning states, which contribute no useful information for the agent, are excluded from these regions and merged into a single state within the agent-winning region, hence leading to a reduced state space. Therefore, later in the composition procedure, we only need to make use of these reduced dfas instead of the complete dfas of the sub-specifications, improving the efficiency of subsequent synthesis steps.
@@
where $ew$ is a special state collapsed from the set of environment-winning states in $\mathcal{G}$, and the transition function $\delta^{\prime}$ is defined as follows:
@@
Next, we show that a strategy that is not an agent-winning in $\mathcal{G}$ is also not an agent-winning strategy in $\textnormal{\sf\small awr}(\mathcal{G})$. When $g$ is not an agent-winning strategy, there exists a play $\tau_{g}$ produced by $g$ that is an environment-winning play. Therefore, either $\tau_{g}$ stays within $\textnormal{\sf\small awin}(\mathcal{G})$ but never visits $F$ or $\tau_{g}$ visits a state $s_{i}\not\in\textnormal{\sf\small awin}(\mathcal{G})$ for some $i\geq 0$. In the former case, Equation (1) assures that $\tau_{g}$ also exists when playing $g$ on $\textnormal{\sf\small awr}(\mathcal{G})$. In this case, $g$ is not an agent-winning strategy on $\textnormal{\sf\small awr}(\mathcal{G})$. In the latter case, there exists a new play $\tau^{\prime}_{g}$ that share the same prefix $h=s_{0}s_{1}\cdots s_{i-1}$ before $\tau_{g}$ visits $s_{i}\not\in\textnormal{\sf\small awin}(\mathcal{G})$ (ii indicates the first occurrence of such state), and $\tau^{\prime}_{g}=h\cdot(ew)^{\omega}$ since all the environment-winning states are collapsed to $ew$. $\tau^{\prime}_{g}$ is indeed an environment-winning play, hence $g$ is not an agent-winning strategy on $\textnormal{\sf\small awr}(\mathcal{G})$. ∎
@@
$g$ is an agent-winning strategy for $\mathcal{G}$ iff every play induced by $g$ terminates at an accepting state of $\mathcal{G}$ which is equivalent to saying that the corresponding traces are accepted by $\mathcal{D}$ and consequently by $\mathcal{G}_{m}$ as $\mathcal{L}(\mathcal{G})=\mathcal{L}(\mathcal{G}_{m})$. Thus, $g$ is also an agent-winning strategy for $\mathcal{G}_{m}$. ∎
@@ 4.2 Individual Composition
The computation of $\textnormal{\sf\small awr}(\mathcal{G}_{\psi_{2}})$ is implemented using the API $\textnormal{{GetAwr}}()$ (Line 2), which adapts existing ltlf synthesis approaches. For a given dfa game, the key to building the agent-winning region is computing the set of the agent-winning states. When applying the on-the-fly synthesis approach to compute the agent-winning region, the forward search must not terminate as soon as the current state is identified as agent-winning. Instead, it should explore all possible agent choices and continue the search recursively. Due to space constraints, the details of computing the agent-winning region on the fly are provided in the technical appendix.
@@ 5 Experimental Evaluation
We implemented the compositional synthesis approach in a prototype tool called Cosy and compared it against state-of-the-art ltlf synthesizers. This section presents experimental results demonstrating that our compositional ltlf synthesis approach outperforms existing ltlf synthesizers. Among the two composition variants, the incremental composition demonstrates better performance than the individual composition. Therefore, unless otherwise specified, the default setting in Cosy is the incremental composition variant. The complete experimental setup, including the source code of Cosy, the benchmarks, the baseline tools, and the original logs from the experiments, is available at [1].

### v2

@@ 1 Introduction
We implemented our synthesis framework in a prototype tool called Cosy. To evaluate the efficiency of the proposed composition techniques, we conducted an empirical evaluation by comparing Cosy with state-of-the-art ltlf synthesizers. Experimental results show that our compositional synthesis framework solves a notable number of instances that other synthesizers cannot handle, demonstrating that it effectively combines the merits of both backward and forward synthesis approaches for ltlf specifications. A detailed comparison of the two composition variants shows that, while the incremental variant solves slightly more instances, both variants have their strengths and advantages. The full version of the paper including appendix is available in [17].
@@ 2 Preliminaries
Definition 1 (Winning Strategy).
Given an ltlf synthesis specification $\left(\varphi,\mathcal{X},\mathcal{Y}\right)$, an agent strategy $g:(2^{\mathcal{X}})^{*}\to 2^{\mathcal{Y}}$ is a winning strategy for $\varphi$ iff for every infinite sequence $\lambda=X_{0}X_{1}\cdots\in(2^{\mathcal{X}})^{\omega}$ of propositional interpretations over $\mathcal{X}$, i.e., every possible environment behaviours, there is $k\geq 0$ such that $\rho\models\varphi$ holds, where $\rho=(X_{0}\cup g(\epsilon))(X_{1}\cup g(X_{0}))\cdots(X_{k}\cup g(X_{0}\cdots X_{k-1}))$
Definition 2 (ltlf Realizability).
Given an ltlf synthesis specification $\left(\varphi,\mathcal{X},\mathcal{Y}\right)$, it is realizable iff there exists a winning strategy. An ltlf specification is unrealizable if it is not realizable. The ltlf realizability problem is to determine whether an ltlf specification $\left(\varphi,\mathcal{X},\mathcal{Y}\right)$ is realizable.
Definition 3 (ltlf Synthesis).
The ltlf synthesis problem (for a realizable specification) is to compute a winning strategy.
@@ 3 Theoretical Foundations
To formulate the composition of strategies, we utilize the agent-winning region in dfa games. Intuitively, the agent-winning region, consisting of all agent-winning states, captures all possible ways the agent can win the game, regardless of how the environment behaves. Consequently, it represents all the information required from each sub-specification to realize the original specification. Notably, the environment-winning states, which contribute no useful information for the agent, are excluded from these regions and merged into a single state, hence leading to a reduced state space. Therefore, later in the composition procedure, we only need to make use of these reduced dfas instead of the complete dfas of the sub-specifications, improving the efficiency of subsequent synthesis steps.
@@
where $ew$ is a special state representing the set of environment-winning states in $\mathcal{G}$. The transition function $\delta^{\prime}$ is defined as follows:
@@
We now show that a strategy $g$ that is not agent-winning in $\mathcal{G}$ is not agent-winning in $\textnormal{\sf\small awr}(\mathcal{G})$. If $g$ is not agent-winning, there exists a play $\tau_{g}$ produced by $g$ that is an environment-winning play. Therefore, either $\tau_{g}$ stays within $\textnormal{\sf\small awin}(\mathcal{G})$ but never visits $F$ or $\tau_{g}$ visits a state $s_{i}\not\in\textnormal{\sf\small awin}(\mathcal{G})$ for some $i\geq 0$. In the former case, Equation (1) assures that $\tau_{g}$ also exists when playing $g$ on $\textnormal{\sf\small awr}(\mathcal{G})$ such that $g$ is not an agent-winning strategy on $\textnormal{\sf\small awr}(\mathcal{G})$. In the latter case, there exists a new play $\tau^{\prime}_{g}$ that shares the same prefix $h=s_{0}s_{1}\cdots s_{i-1}$ before $\tau_{g}$ visits $s_{i}\not\in\textnormal{\sf\small awin}(\mathcal{G})$ (ii indicates the first occurrence of such state), and $\tau^{\prime}_{g}=h\cdot(ew)^{\omega}$ (note that $ew$ represents all the environment-winning states). $\tau^{\prime}_{g}$ is indeed an environment-winning play, hence $g$ is not an agent-winning strategy on $\textnormal{\sf\small awr}(\mathcal{G})$. ∎
@@
$g$ is an agent-winning strategy for $\mathcal{G}$ iff every play induced by $g$ terminates at an accepting state of $\mathcal{G}$, hence the corresponding traces are accepted by $\mathcal{D}$ and consequently by $\mathcal{G}_{m}$ as $\mathcal{L}(\mathcal{G})=\mathcal{L}(\mathcal{G}_{m})$. Therefore, $g$ is also an agent-winning strategy for $\mathcal{G}_{m}$. ∎
@@ 4.2 Individual Composition
The computation of $\textnormal{\sf\small awr}(\mathcal{G}_{\psi_{2}})$ is implemented using the API $\textnormal{{GetAwr}}()$ (Line 2), which adapts existing ltlf synthesis approaches. For a given dfa game, the key to building the agent-winning region is computing the set of agent-winning states. When applying the on-the-fly synthesis approach to compute the agent-winning region, the forward search must not terminate as soon as the current state is identified as agent-winning. Instead, it should explore all possible agent choices and continue the search recursively.
@@ 5 Experimental Evaluation
We implemented the compositional synthesis approach in a prototype tool called Cosy [1] and compared it against state-of-the-art ltlf synthesizers. This section presents experimental results demonstrating that our compositional ltlf synthesis approach outperforms existing ltlf synthesizers. Among the two composition variants, the incremental composition demonstrates better performance than the individual composition. Therefore, unless otherwise specified, the default setting in Cosy is the incremental composition variant.
@@ 6 Concluding Remarks
{ack}
We sincerely thank the anonymous reviewers for their valuable comments, which significantly improved the quality of this paper. This work was supported by the National Natural Science Foundation of China (Grants #62372178 and #U21B2015) and the Shanghai Collaborative Innovation Center of Trusted Industry Internet Software.